OpenStudy (anonymous):

The probability of drawing a dime from a bag of coins is 7/10 . What is the probability of drawing a coin that is not a dime from the bag?

OpenStudy (kropot72):

P(not a dime) = 1 - 7/10 = 3/10 or 0.3